# Jonas E. Arias Federal Reserve Bank of Philadelphia Juan Rubio-Ramírez

### v1

@@
(May 29, 2025)
We develop a new algorithm for inference based on SVARs identified with sign restrictions. The key insight of our algorithm is to break apart from the accept-reject tradition associated with sign-identified SVARs. We show that embedding an elliptical slice sampling within a Gibbs sampler approach can deliver dramatic gains in speed and turn previously infeasible applications into feasible ones. We provide a tractable example to illustrate the power of the elliptical slice sampling applied to sign-identified SVARs. We demonstrate the usefulness of our algorithm by applying it to a well-known small-SVAR model of the oil market featuring a tight identified set as well as to large SVAR model with more than 100 sign restrictions.
@@ 1 Introduction
The conventional method of Bayesian inference in SVARs identified with sign restrictions—proposed by Faust (1998); Canova and De Nicoló (2002); Uhlig (2005) and extended by Rubio-Ramírez, Waggoner, and Zha (2010)—is one of the most popular tools available for assessing the dynamic causal effects of structural shocks in macroeconomics. To date, thousands of studies have adopted this standard accept-reject algorithm. Despite recent advances (Chan, Matthes, and Yu, 2025, have proposed an improved accept-reject algorithm by exploiting permutations and sign alternation of the standard draw from the uniform prior density over the set of orthogonal matrices), the big data era has strained the capacity of this class of algorithms to the point that in some cases it becomes infeasible as identified sets become tight. Similar challenges arise even in lower-dimensional SVARs when incorporating rank restrictions, such as elasticity bounds, as discussed in Kilian and Murphy (2012); Amir-Ahmadi and Drautzburg (2021); these ranking restrictions can meaningfully sharpen inference at the cost of tight identified sets. The adoption of large-scale SVARs has become increasingly common in macroeconomic research following the foundational work of Leeper, Sims, and Zha (1996), who introduced medium-sized systems to analyze monetary policy shocks. This shift toward higher-dimensional models was significantly influenced by Bańbura, Giannone, and Reichlin (2010), who highlighted the advantages of including a broad set of variables for improving both predictive accuracy and structural interpretation. Subsequent contributions have demonstrated the practicality and value of large SVARs in various contexts, including the studies by Carriero, Kapetanios, and Marcellino (2009); Koop (2013); Ellahie and Ricco (2017); Crump, Eusepi, Giannone, Qian, and Sbordone (2021, 2025), which illustrate how such models can be effectively utilized in empirical macroeconomic analysis.
@@
In this paper, we break apart with accept-reject tradition and show that embedding an elliptical slice sampling within a Gibbs sampler approach can deliver dramatic gains in speed and turn previously infeasible applications into feasible ones. The objective is to obtain draws from the posterior distribution of the orthogonal reduced-form parameters conditional on the sign restrictions. To accomplish such a goal, the conventional approach draws from the posterior distribution of the orthogonal reduced-form parameters and uses an accept-reject algorithm to discard the draws that do not satisfy the restrictions. Consequently, as more restrictions are incorporated to sharpen inference, this approach is destined to fail. The innovations in Chan, Matthes, and Yu (2025) will certainly expand the set of models that researchers can consider. Notwithstanding, it will eventually fail as the sets become even tighter. In contrast, we use a Gibbs sampler to iteratively draw from the posterior distributions conditional on the sign restrictions, making the accept-reject step unnecessary.
@@
We illustrate our algorithm with two applications. In the first application, we replicate Kilian and Murphy (2014), a model of the world oil market where the standard application of the accept-reject algorithm fails. To get around this infeasibility, Kilian and Murphy (2014) consider an approach similar to the one in Chan, Matthes, and Yu (2025) by exploiting permutations and sign alternation. As we will show below, our algorithm can efficiently handle this application. In the second application, we re-visit the structural analysis in Chan, Matthes, and Yu (2025) who use Crump et al.’s (2025) large SVAR model of the U.S. economy to identify 8 structural shocks. We show that as we augment the number of shocks under analysis, the efficient accept-reject proposed by Crump, Eusepi, Giannone, Qian, and Sbordone (2025) slows down noticeable until it becomes impractical. In contrast, our the computational time of our Gibbs sampler is largely insensitive of the number of identified structural shocks.
@@ 2 The Model
where ${\bf{B}}={\bf{A}}_{+}{\bf{A}}_{0}^{-1}$, ${\bf{u}}_{t}^{\prime}={\bm{\varepsilon}}_{t}^{\prime}{\bf{A}}_{0}^{-1}$, and $\mathbb{E}\left[{\bf{u}}_{t}{\bf{u}}_{t}^{\prime}\right]={\bf{\Sigma}}=\left({\bf{A}}_{0}{\bf{A}}_{0}^{\prime}\right)^{-1}$. The matrices ${\bf{B}}$ and ${\bf{\Sigma}}$ are the reduced-form parameters, while ${\bf{A}}_{0}$ and ${\bf{A}}_{+}$ are the structural parameters. While $\operatorname{\mathbf{B}}$ is a $m\times n$ matrix, $\operatorname{\mathbf{\Sigma}}$ belong to the set $\mathcal{S}(n)$, which is the set of $n\times n$ positive definite matrices.
@@ 2.3 Inference Based on Sign Restrictions
While this algorithm has been widely adopted, it is well known that there are cases in which the identified set is narrow limiting the efficiency of the algorithm (see e.g. Kilian and Murphy, 2014; Baumeister and Hamilton, 2024; Chan, Matthes, and Yu, 2025; Read and Zhu, 2025). In the next section, we use a simple example to show it shortcomings. We will also show how a carefully designed ESS algorithm is not subject to this limitation and delivers dramatic speed gains. Importantly, Chan, Matthes, and Yu (2025) show a new numerically efficient version of Algorithm 1 that facilitates the drawing for a large number of structural restrictions. When comparing our algorithm to the traditional accept-reject approach we will use this efficient version as the benchmark.
@@ 3 The Problem with Accept-Reject Sampling
${\bf{y}}_{t}^{\prime}=\left(y_{t,1},y_{t,2}\right)={\bm{\varepsilon}}_{t}^{\prime}({\bf\Sigma}_{tr}{\bf Q})^{\prime}$
@@
Given the above, it is easy to see that the impact of the first shock on $y_{t,1}$ and $y_{t,2}$ can be written as $\ell_{11}={q}_{11}$ and $\ell_{21}=-0.9{q}_{11}+{q}_{21}$, where $\ell_{i1}$ and $q_{i1}$ are the $i$-th row and $1$-th column entry of ${\bf L}_{0}$ and ${\bf Q}$, respectively. Let us now impose sign restrictions requiring that $\ell_{11}$ and $\ell_{21}$ are nonnegative. The imposed sign restrictions imply $q_{11}\geq 0$ and $q_{21}\geq 0.9q_{11}$. Figure 1(a) illustrates this setup graphically. The green circle represents the domain of ${\bf q}_{1}=\left(q_{11},q_{21}\right)^{\prime}$, while the red arc highlights the identified set that satisfies the imposed sign restrictions.
@@
The efficiency of the posterior simulator based on this type of acceptance-rejection algorithm depends heavily on the size of the identified set. As the identified set becomes tighter, we naturally expect to discard a larger number of draws. Indeed, the expected number of draws required to satisfy the sign restrictions is inversely proportional to the probability of meeting those restrictions. Figure 1(b) illustrates this relationship: the green line plots the expected number of draws needed to satisfy the sign restrictions as a function of the size of the identified set (i.e., the length of the red arc). More specifically, we generate smaller identified sets by gradually moving the left endpoint of the red arc toward its right endpoint. As shown in the figure, the expected number of draws required to meet the sign restrictions increases hyperbolically as the identified set shrinks. In realistic scenarios, as illustrated later in our empirical applications, the number of draws required can become quite large rendering the algorithm inefficient.
@@
$\bf{x}_{1}^{(\star)}=\bm{\nu}\sin(\theta)+\bf{x}_{1}^{(0)}\cos(\theta)\text{ with }\theta\in\left[0,2\pi\right]$
@@
The fact that the candidate set for $\theta$ shrinks exponentially is an appealing feature, as it significantly reduces the number of candidate draws $\bf{q}_{1}$ needed to satisfy the restrictions. This efficiency gain becomes particularly important when the dimension of the model increases since generating new draws of $\bf{q}_{1}$ is computationally costly. Figure 1(b) (red line) displays the average number of trials required by ESS to generate an accepted draw of $\bf{q}_{1}$ within the identified set as a function of the length of the identified set. The number of required trials using ESS grows only logarithmically, whereas the conventional acceptance-rejection sampler grows hyperbolically. [^1] To make a fair comparison, one should account for the serial correlation introduced by the Gibbs sampling algorithm, since the accept/reject algorithm generates i.i.d. draws. As demonstrated in a later section, we compute the effective sample size and find that the number of draws required to obtain one i.i.d.-equivalent sample ranges from 1.04 to 1.35. Therefore, this adjustment does not alter the main conclusion illustrated in the figure.
@@ 4 An Algorithm
The first equality follows from Bayes rule. The second equality is straightforward because $NIW_{(\tilde{\nu},\tilde{\bf\Phi},\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}})$ does not depend on $\operatorname{\mathbf{Q}}$. The proportionality follows from Assumption 1.1.
@@
The first equality follows from Bayes rule and the following factorization of the Normal-Inverse Wishart density: $NIW_{(\tilde{\nu},\tilde{\bf\Phi},\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}})$ as $N_{(\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})IW_{(\tilde{\nu},\tilde{\bf\Phi})}(\operatorname{\mathbf{\Sigma}})$. As we demonstrate below, this factorization is convenient to demonstrate that the posterior under analysis can be induced by a kernel that contains a Gaussian density among its components. The proportionality follows from Assumption 1.2.
Let us now explain how to draw from $p(\operatorname{\mathbf{\Sigma}}\mid\operatorname{\mathbf{B}},\operatorname{\mathbf{Q}},(\mathbf{y}_{t})_{t=1}^{T},\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0)$. We will exploit the mapping between an $n\times\tilde{\nu}$ matrix $\operatorname{\mathbf{R}}\sim N({\bf 0},\tilde{\operatorname{\mathbf{\Phi}}}^{-1})$ and $\operatorname{\mathbf{\Sigma}}$. More specifically, if $\operatorname{\mathbf{\Sigma}}=(\operatorname{\mathbf{R}}\operatorname{\mathbf{R}}^{\prime})^{-1}$ we have that $\operatorname{\mathbf{\Sigma}}\sim IW(\tilde{\nu},\tilde{\operatorname{\mathbf{\Phi}}})$. Let us denote such a mapping by $\operatorname{\mathbf{\Sigma}}=\varsigma(\operatorname{\mathbf{R}})$. Hence, if one draws $\operatorname{\mathbf{R}}$ from $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0]N_{(\tilde{\bf\Psi},\varsigma(\operatorname{\mathbf{R}})\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})N_{(\bf{0},\tilde{\bf\Phi}^{-1})}(\operatorname{\mathbf{R}})$ and then transforms the draws using $\operatorname{\mathbf{\Sigma}}=\varsigma(\operatorname{\mathbf{R}})$, it obtain draws from the desired distribution $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0]N_{(\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})IW_{(\tilde{\nu},\tilde{\bf\Phi})}(\operatorname{\mathbf{\Sigma}})$. As before, because $\operatorname{\mathbf{R}}$ is normal, we can sample from $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}}^{i})>0]N_{(\tilde{\bf\Psi},\varsigma(\operatorname{\mathbf{R}})\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})N_{(\bf{0},\tilde{\bf\Phi}^{-1})}(\operatorname{\mathbf{R}})$ using the ESS.
@@
The third and last conditional posterior corresponds to we the posterior of $\operatorname{\mathbf{B}}$ given $(\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})$ and the sign restrictions. In this case, we use Equation (7) to obtain:
@@
The proportionality comes from from Assumption 1.3. Notice that in this case, it is immediately clear that the conditional posterior of interest contains a Gaussian kernel. Hence, we can draw from $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0]N_{(\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})$ using the ESS.

### v2

@@
(May 30, 2025)
We develop a new algorithm for inference based on structural vector autoregressions (SVARs) identified with sign restrictions. The key insight of our algorithm is to break apart from the accept-reject tradition associated with sign-identified SVARs. We show that embedding an elliptical slice sampling within a Gibbs sampler approach can deliver dramatic gains in speed and turn previously infeasible applications into feasible ones. We provide a tractable example to illustrate the power of the elliptical slice sampling applied to sign-identified SVARs. We demonstrate the usefulness of our algorithm by applying it to a well-known small-SVAR model of the oil market featuring a tight identified set, as well as to a large SVAR model with more than 100 sign restrictions.
@@ 1 Introduction
The conventional method of Bayesian inference in SVARs identified with sign restrictions—proposed by Faust (1998); Canova and De Nicoló (2002); Uhlig (2005) and extended by Rubio-Ramírez, Waggoner, and Zha (2010)—is one of the most popular tools available for assessing the dynamic causal effects of structural shocks in macroeconomics. To date, thousands of studies have adopted this standard accept-reject algorithm. Despite recent advances—Chan, Matthes, and Yu (2025) have proposed an improved accept-reject algorithm by exploiting permutations and sign alternation of the standard draw from the uniform prior density over the set of orthogonal matrices—the big data era has strained the capacity of this class of algorithms to the point that in some cases it becomes infeasible as identified sets become tight. Similar challenges arise even in lower-dimensional SVARs when incorporating rank restrictions, such as elasticity bounds, as discussed in Kilian and Murphy (2012); Amir-Ahmadi and Drautzburg (2021); these ranking restrictions can meaningfully sharpen inference at the cost of tight identified sets. The adoption of large-scale SVARs has become increasingly common in macroeconomic research following the foundational work of Leeper, Sims, and Zha (1996), who introduced medium-sized systems to analyze monetary policy shocks. This shift toward higher-dimensional models was significantly influenced by Bańbura, Giannone, and Reichlin (2010), who highlighted the advantages of including a broad set of variables for improving both predictive accuracy and structural interpretation. Subsequent contributions have demonstrated the practicality and value of large SVARs in various contexts, including the studies by Carriero, Kapetanios, and Marcellino (2009); Koop (2013); Ellahie and Ricco (2017); Crump, Eusepi, Giannone, Qian, and Sbordone (2021, 2025), which illustrate how such models can be effectively utilized in empirical macroeconomic analysis.
@@
In this paper, we break with the accept-reject tradition and show that embedding an elliptical slice sampling within a Gibbs sampler approach can deliver dramatic gains in speed and turn previously infeasible applications into feasible ones. The objective is to obtain draws from the posterior distribution of the orthogonal reduced-form parameters conditional on the sign restrictions. To accomplish such a goal, the conventional approach draws from the posterior distribution of the orthogonal reduced-form parameters and uses an accept-reject algorithm to discard the draws that do not satisfy the restrictions. Consequently, as more restrictions are incorporated to sharpen inference, this approach is destined to fail. The innovations in Chan, Matthes, and Yu (2025) will certainly expand the set of models that researchers can consider. Notwithstanding, it will eventually fail as the sets become even tighter. In contrast, we use a Gibbs sampler to iteratively draw from the posterior distributions conditional on the sign restrictions, making the accept-reject step unnecessary.
@@
We illustrate our algorithm with two applications. In the first application, we replicate Kilian and Murphy (2014), a model of the world oil market where the standard application of the accept-reject algorithm fails. To get around this infeasibility, Kilian and Murphy (2014) consider an approach similar to the one in Chan, Matthes, and Yu (2025) by exploiting permutations and sign alternation. As we will show below, our algorithm can efficiently handle this application. In the second application, we re-visit the structural analysis in Chan, Matthes, and Yu (2025) who use Crump et al.’s (2025) large SVAR model of the U.S. economy to identify 8 structural shocks. We show that as we augment the number of shocks under analysis, the efficient accept-reject proposed by Crump, Eusepi, Giannone, Qian, and Sbordone (2025) slows down noticeably until it becomes impractical. In contrast, the computational time of our Gibbs sampler is largely insensitive of the number of identified structural shocks.
@@ 2 The Model
where ${\bf{B}}={\bf{A}}_{+}{\bf{A}}_{0}^{-1}$, ${\bf{u}}_{t}^{\prime}={\bm{\varepsilon}}_{t}^{\prime}{\bf{A}}_{0}^{-1}$, and $\mathbb{E}\left[{\bf{u}}_{t}{\bf{u}}_{t}^{\prime}\right]={\bf{\Sigma}}=\left({\bf{A}}_{0}{\bf{A}}_{0}^{\prime}\right)^{-1}$. The matrices ${\bf{B}}$ and ${\bf{\Sigma}}$ are the reduced-form parameters, while ${\bf{A}}_{0}$ and ${\bf{A}}_{+}$ are the structural parameters. While $\operatorname{\mathbf{B}}$ is a $m\times n$ matrix, $\operatorname{\mathbf{\Sigma}}$ belongs to the set $\mathcal{S}(n)$, which is the set of $n\times n$ positive definite matrices.
@@ 2.3 Inference Based on Sign Restrictions
While this algorithm has been widely adopted, it is well known that there are cases in which the identified set is narrow limiting the efficiency of the algorithm (see e.g. Kilian and Murphy, 2014; Baumeister and Hamilton, 2024; Chan, Matthes, and Yu, 2025; Read and Zhu, 2025). In the next section, we use a simple example to show its shortcomings. We will also show how a carefully designed ESS algorithm is not subject to this limitation and delivers dramatic speed gains. Importantly, Chan, Matthes, and Yu (2025) show a new numerically efficient version of Algorithm 1 that facilitates the drawing for a large number of structural restrictions. When comparing our algorithm to the traditional accept-reject approach we will use this efficient version as the benchmark.
@@ 3 The Problem with Accept-Reject Sampling
${\bf{y}}_{t}^{\prime}=\left(y_{t,1},y_{t,2}\right)={\bm{\varepsilon}}_{t}^{\prime}({\bf\Sigma}_{tr}{\bf Q})^{\prime},$
@@
Given the above, it is easy to see that the impact of the first shock on $y_{t,1}$ and $y_{t,2}$ can be written as $\ell_{11}={q}_{11}$ and $\ell_{21}=-0.9{q}_{11}+{q}_{21}$, where $\ell_{i1}$ and $q_{i1}$ are the $(i,1)$ entries of ${\bf L}_{0}$ and ${\bf Q}$, respectively. Let us now impose sign restrictions requiring that $\ell_{11}$ and $\ell_{21}$ are nonnegative. The imposed sign restrictions imply $q_{11}\geq 0$ and $q_{21}\geq 0.9q_{11}$. Figure 1(a) illustrates this setup graphically. The green circle represents the domain of ${\bf q}_{1}=\left(q_{11},q_{21}\right)^{\prime}$, while the red arc highlights the identified set that satisfies the imposed sign restrictions.
@@
The efficiency of the posterior simulator based on this type of acceptance-rejection algorithm depends heavily on the size of the identified set. As the identified set becomes tighter, we naturally expect to discard a larger number of draws. Indeed, the expected number of draws required to satisfy the sign restrictions is inversely proportional to the probability of meeting those restrictions. Figure 1(b) illustrates this relationship: the green line plots the expected number of draws needed to satisfy the sign restrictions as a function of the size of the identified set (i.e., the length of the red arc). More specifically, we generate smaller identified sets by gradually moving the left endpoint of the red arc toward its right endpoint. As shown in the figure, the expected number of draws required to meet the sign restrictions increases hyperbolically as the identified set shrinks. In realistic scenarios, as illustrated later in our empirical applications, the number of draws required can become quite large, rendering the algorithm inefficient.
@@
$\bf{x}_{1}^{(\star)}=\bm{\nu}\sin(\theta)+\bf{x}_{1}^{(0)}\cos(\theta)\text{ with }\theta\in\left[0,2\pi\right],$
@@
The fact that the candidate set for $\theta$ shrinks exponentially is an appealing feature, as it significantly reduces the number of candidate draws $\bf{q}_{1}$ needed to satisfy the restrictions. This efficiency gain becomes particularly important when the dimension of the model increases since generating new draws of $\bf{q}_{1}$ is computationally costly. Figure 1(b) (red line) displays the average number of trials required by ESS to generate an accepted draw of $\bf{q}_{1}$ within the identified set as a function of the length of the identified set. The number of required trials using ESS grows only logarithmically, whereas the conventional acceptance-rejection sampler grows hyperbolically. [^1] To make a fair comparison, one should account for the serial correlation introduced by the Gibbs sampling algorithm, since the accept-reject algorithm generates i.i.d. draws. As demonstrated in a later section, we compute the effective sample size and find that the number of draws required to obtain one i.i.d.-equivalent sample ranges from 1.04 to 1.35. Therefore, this adjustment does not alter the main conclusion illustrated in the figure.
@@ 4 An Algorithm
The first equality follows from Bayes’ rule. The second equality is straightforward because $NIW_{(\tilde{\nu},\tilde{\bf\Phi},\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}})$ does not depend on $\operatorname{\mathbf{Q}}$. The proportionality follows from Assumption 1.1.
@@
The first equality follows from Bayes’ rule and the following factorization of the normal-inverse Wishart density: $NIW_{(\tilde{\nu},\tilde{\bf\Phi},\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}})$ as $N_{(\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})IW_{(\tilde{\nu},\tilde{\bf\Phi})}(\operatorname{\mathbf{\Sigma}})$. As we demonstrate below, this factorization is convenient to demonstrate that the posterior under analysis can be induced by a kernel that contains a Gaussian density among its components. The proportionality follows from Assumption 1.2.
Let us now explain how to draw from $p(\operatorname{\mathbf{\Sigma}}\mid\operatorname{\mathbf{B}},\operatorname{\mathbf{Q}},(\mathbf{y}_{t})_{t=1}^{T},\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0)$. We will exploit the mapping between an $n\times\tilde{\nu}$ matrix $\operatorname{\mathbf{R}}\sim N({\bf 0},\tilde{\operatorname{\mathbf{\Phi}}}^{-1})$ and $\operatorname{\mathbf{\Sigma}}$. More specifically, if $\operatorname{\mathbf{\Sigma}}=(\operatorname{\mathbf{R}}\operatorname{\mathbf{R}}^{\prime})^{-1}$ we have that $\operatorname{\mathbf{\Sigma}}\sim IW(\tilde{\nu},\tilde{\operatorname{\mathbf{\Phi}}})$. Let us denote such a mapping by $\operatorname{\mathbf{\Sigma}}=\varsigma(\operatorname{\mathbf{R}})$. Hence, if one draws $\operatorname{\mathbf{R}}$ from $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0]N_{(\tilde{\bf\Psi},\varsigma(\operatorname{\mathbf{R}})\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})N_{(\bf{0},\tilde{\bf\Phi}^{-1})}(\operatorname{\mathbf{R}})$ and then transforms the draws using $\operatorname{\mathbf{\Sigma}}=\varsigma(\operatorname{\mathbf{R}})$, one obtains draws from the desired distribution $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0]N_{(\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})IW_{(\tilde{\nu},\tilde{\bf\Phi})}(\operatorname{\mathbf{\Sigma}})$. As before, because $\operatorname{\mathbf{R}}$ is normal, we can sample from $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}}^{i})>0]N_{(\tilde{\bf\Psi},\varsigma(\operatorname{\mathbf{R}})\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})N_{(\bf{0},\tilde{\bf\Phi}^{-1})}(\operatorname{\mathbf{R}})$ using the ESS.
@@
The third and last conditional posterior corresponds to the posterior of $\operatorname{\mathbf{B}}$ given $(\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})$ and the sign restrictions. In this case, we use Equation (7) to obtain:
@@
The proportionality comes from Assumption 1.3. Notice that in this case, it is immediately clear that the conditional posterior of interest contains a Gaussian kernel. Hence, we can draw from $[\bm{S}(\operatorname{\mathbf{B}},\operatorname{\mathbf{\Sigma}},\operatorname{\mathbf{Q}})>0]N_{(\tilde{\bf\Psi},\operatorname{\mathbf{\Sigma}}\otimes\tilde{\bf\Omega})}(\operatorname{\mathbf{B}})$ using the ESS.